a. Suppose the number of employed people in Pakistan is 20.5 million. The unemployment rate in this economy is 8.8 percent, and the labor force participation rate is 65 percent. (1.5 Marks)
i. What is the size of the labor force and working-age population?
ii. How many people are unemployed?
1
Expert's answer
2020-11-27T12:46:50-0500

i. The number of employed are 20.5 million.

Also we are given the unemployment rate as 8.8 %.

This means that employment rate is 100 – 8.8 % = 91.2 %

From this the total labor force could be determined:

Total labor force "= \frac{20.5 \times 100}{91.2} = 22.47 \;million"

So, labor force participation is $22.47 million.

Now the labor force is 65 % of total adult population.

Total adult population "= \frac{22.47 \times 100}{65}= 34.58 \;million"

ii. Number of unemployed "= \frac{8.8 \times 22.47}{100} =1.97 \;million"